A particular paint has a density of 0.914g/cm3. You need to cover a wall that is 7.6m long and 2.74m high with a paint layer 0.13mm thick. What volume of paint (in litres) is required? What is the mass (in grams) of the paint layer?

1
Expert's answer
2022-05-17T15:14:04-0400

Solution:

Density of paint = 0.914 g cm−3 = 914 kg m−3


Volume of paint = Long × High × Thick

Volume of paint = (7.6 m) × (2.74 m) × (0.00013 m) = 0.002707 m3

Volume of paint = (0.002707 m3) × (1000 L / 1 m3) = 2.707 L

Volume of paint = 2.707 L


Density = Mass / Volume

Therefore,

Mass of paint layer = Density of paint × Volume of paint

Mass of paint layer = (914 kg m−3) × (0.002707 m3) = 2.4742 kg

Mass of paint layer = (2.4742 kg) × (1000 g / 1 kg) = 2474.2 g

Mass of paint layer = 2474.2 g


Answer:

The volume of paint is 2.707 L

The mass of the paint layer is 2474.2 grams